OpenStudy (anonymous):

May you help me on this problem, please? I know it is not 2:1 . @ganeshie8 @jhonyy9 @surjithayer

OpenStudy (anonymous):

calculate the volume using the formula v=(4/3)πr² for each then compare to find the ratio

OpenStudy (anonymous):

4! :D I really thought it was 4:1 !! :DD

OpenStudy (anonymous):

\[r _{1}=\frac{ 18 }{ 2 }=9 inch,r _{2}=\frac{ 9 }{ 2 } inch\] \[\frac{ V _{1} }{ V _{2} }=\frac{ \frac{ 1 }{ 3 } \pi r _{1}^3}{ \frac{ 1 }{ 3 }\pi r _{2} ^3}=\frac{ r _{1} ^3}{ r _{2} ^3}\]

jhonyy9 (jhonyy9):

so in case of hat A r= 9 in case of hate B r=4,5 you know area =pir^2 hat A has area 81pi and hat B area 18,5pi ratio of the volum supposed that are the same height will be ratio 81/18,5 = ?

OpenStudy (anonymous):

18.5 you mean? If so, that would =4.378

OpenStudy (anonymous):

\[\frac{ V _{1} }{ V _{2} }=\frac{ 9^3 }{ \left( \frac{ 9 }{ 2 } \right)^3 }=\frac{ 9^3 }{ 9^3 }\times2^3=\frac{ 8 }{ 1 }\]

jhonyy9 (jhonyy9):

4,5 squared = 20,25 will be right ---- sorry

OpenStudy (anonymous):

so \[V _{1}:V _{2}=8:1\]

jhonyy9 (jhonyy9):

so my opinion is 4:1

OpenStudy (anonymous):

What!?? Why are you putting commas? Also, why is everyone getting different answers!?? 4:1, 8:1, 2 .. @ganeshie8 I need your help!

OpenStudy (anonymous):